OpenStudy (anonymous):

A ball hangs on a string and is rotated in a circle(m=.275, r=.850). If the String breaks when its tension exceeds 22.5N, What is the maximum speed the ball can have at the bottom before it breaks?

OpenStudy (anonymous):

Vertical or horizontal circle. Like a water wheel on an old corn mill or like a record on a turn table. (These are probably terrible analogies for the younger crowd.)

OpenStudy (anonymous):

Vertical circles like the water wheel

OpenStudy (anonymous):

Since we only care about the case when the mass is at the bottom of the circle, this simplifies things. Draw a FBD. We assume \(\omega\) (angular velocity) is constant. Let's balance the forces. \[\sum F = ma = 0 = T - mg\]T is equal to the centripetal force. \[T = {mv^2 \over r}\] With me?

OpenStudy (anonymous):

Yes I am with you so far

OpenStudy (anonymous):

Hmm. We are over defined.

OpenStudy (anonymous):

So I would set that up as \[22.5N = \frac{mv^2}{r}-mg\] When I do that I get a different answer than the book says

OpenStudy (anonymous):

That's because I did it wrong. T doesn't equal the centripetal force. \[\sum F_r = ma_c = T - mg\] Now, \[ma_c = {mv^2 \over r}\] This yields, \[{mv^2 \over r} = 22.5 - mg\]

OpenStudy (anonymous):

Awesome bro Thanks a ton.

OpenStudy (anonymous):

I attempted to set it up the same way you did the first time.

OpenStudy (anonymous):

Yeah. Sorry about the misstep. Hopefully you're clear on the "why" behind the proper set-up.

OpenStudy (anonymous):

If not, let me know and I'll lay it out more fundamentally.

OpenStudy (anonymous):

No i'm with you bur for whatever reason I wanted to set ma=22.5 instead of setting tension to 22.5. Thanks for the help

OpenStudy (anonymous):

Great. Good luck.